# A Regret Perspective on Online Selective Generation

### v1

@@ 4 Method: Online Selective Generation with Partial Feedback Unlocking
Notations. We first introduce our choice of the selection function and simplified notations. As in conventional selective prediction literature, we consider the scalar parameterization of a selection function given a scoring function $f:\mathcal{{X}}\times\mathcal{{Y}}\to{[0,1)}$, i.e., $\hat{{s}}(\mathbf{x}_{t},G(\mathbf{x}_{t}))\coloneqq\mathbbm{1}\left(f(\mathbf{x}_{t},G(\mathbf{x}_{t}))\geq\tau\right),$ thus the selective generator $\hat{{S}}$ is parameterized by $\tau$, denoted by $\hat{{S}}(\cdot;\tau)$. Here, the scoring function $f$ measures the likelihood of $G(\mathbf{x}_{t})$ for being an answer of $\mathbf{x}_{t}$. Also, we specifically consider that $\tau$ is chosen from the finely-quantized, finite space of $[0,1]$, denoted by $\mathcal{{H}}$ (i.e., $\tau\in\mathcal{{H}}$). Given this parameterization, we consider the following shorthand to simplify key notations: $a_{t}(\tau)\coloneqq\mathbbm{1}(\hat{{S}}(\mathbf{x}_{t};\tau)=\texttt{IDK})$, which measures the selection inefficiency at time $t$, called inefficiency loss, and $d_{t}(\tau,\alpha)\coloneqq\mathbbm{1}(\hat{{S}}(\mathbf{x}_{t};\tau)\neq\texttt{IDK}\wedge\hat{{S}}(\mathbf{x}_{t};\tau)\neq_{E}\mathbf{y}_{t})-\alpha\mathbbm{1}(\hat{{S}}(\mathbf{x}_{t};\tau)\neq\texttt{IDK})+\alpha$, which measures the violation of the FDR constraint at time $t$ with a margin by $\alpha$ to penalize the IDK response, called FDR loss with a margin. Based on these short hands, we define loss for online learning and bandits that simultaneously considers the FDR constraint and selection inefficiency as follows:
$\ell_{t}(\tau,\alpha)\coloneqq a_{t}(\tau)+\lambda d_{t}(\tau,\alpha)\in\{0,\lambda,1+\lambda\alpha\},$ (6)
where $\lambda$ is a regularized parameter, which will be determined via theoretical analysis.
@@ 4.1 Reduction: From Online Selective Generation to Adversarial Bandits
models selective generators $\mathcal{{H}}$ finite arms $\mathcal{{H}}$
feedback $\mathbbm{1}(\hat{{S}}(\mathbf{x}_{t};\tau_{t})\neq_{E}\mathbf{y}_{t})$ $\ell_{t}(\tau_{t},\alpha)$
@@
Here, we map the components of online selective generation with partial feedback to those of adversarial bandits to leverage existing algorithms and regret bounds. In particular, each parameter $\tau\in\mathcal{{H}}$ of a selective generator is considered as an arm $\tau\in\mathcal{{H}}$ in adversarial bandits. We learn selective generators from partial feedback $\mathbbm{1}(\hat{{S}}(\mathbf{x}_{t};\tau_{t})\neq_{E}\mathbf{y}_{t})$, i.e., whether an output of a selective generator $\hat{{S}}$ with a chosen parameter $\tau_{t}$ is incorrect or not, where $\hat{{S}}(\mathbf{x}_{t};\tau_{t})=\texttt{IDK}$ incurs non-zero feedback by the definition. In adversarial bandits, we consider two-step feedback: (1) get feedback from the nature $\mathbbm{1}(\hat{{S}}(\mathbf{x}_{t};\tau_{t})\neq_{E}\mathbf{y}_{t})$ and (2) use this to compute the loss (6) by the chosen arm $\tau_{t}$. Based on observed loss up to the time horizon $T$, the adversarial bandit learner updates a strategy $p_{t}$ over time to minimize the regret $\textbf{Reg}_{T}$. Then, by leveraging our conversion lemma (Lemma 1), the algorithm provides an FDR bound. See Table 1 for a summary of reduction.
4.2 Algorithm: Exp3 for Online Selective Generation with Partial Feedback Unlocking
@@
As for the adversarial bandit learner, we leverage Exp3 auer2002nonstochastic that is a direct extension of the EW littlestone1994weighted for partial feedback. The key challenge in moving from full to partial feedback is the slow convergence of learning due to the lack of feedback information. To address this, we exploit the unique parameter structure of selective generators to unlock feedback of other arms by a chosen arm, i.e., feedback unlocking. In particular, by the monotonicity of a selection function in $\tau$, i.e., $\hat{{s}}(\mathbf{x})\coloneqq\mathbbm{1}(f_{t}\geq\tau)$ where $f_{t}\coloneqq f(\mathbf{x}_{t},G(\mathbf{x}_{t}))$ and the definition of a selective generator $\hat{{S}}$, the following relation among feedback holds, where $b_{t}\coloneqq\mathbbm{1}(\hat{{S}}(\mathbf{x}_{t};\tau_{t})\neq_{E}\mathbf{y}_{t})$ is observed feedback by a chosen arm $\tau_{t}$:
[…]
@@
To exploit this feedback unlocking within Exp3, we propose a novel unbiased loss estimator
@@
$\ell_{t}(\tau,\alpha\mid\mathcal{{H}}_{t}({\tau_{t}}))\coloneqq\frac{\ell_{t}(\tau,\alpha)}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\bar{\tau}}))\cdot p_{t}(\bar{\tau})}\cdot\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\tau_{t}}))\in[0,\infty),$ (8)
@@
where $\mathcal{{H}}_{t}({\tau_{t}})\coloneqq\{\tau\in\mathcal{{H}}\mid\tau\leq f_{t}\}$ if $\hat{{S}}(\mathbf{x}_{t};\tau_{t})\neq\texttt{IDK}$ and $\mathcal{{H}}_{t}(\tau_{t})\coloneqq\{\tau\in\mathcal{{H}}\mid\tau>f_{t}\}$ otherwise. Note that $\tau\in\mathcal{{H}}_{t}(\tau)$ by its definition and it is an unbiased estimator, i.e., $\mathbbm{E}_{\tau_{t}\sim p_{t}}\ell_{t}(\tau,\alpha\mid\mathcal{{H}}_{t}({\tau_{t}}))=\ell_{t}(\tau,\alpha)$. See (29) in Appendix G.4 for a proof on the unbiasedness and Algorithm 1 for our extension of Exp3 with partial feedback unlocking.
Connection to Exp3. Our algorithm can be reduced to Exp3 by letting $\mathcal{{H}}_{t}({\tau})=\{\tau_{t}\}$ for any $\tau$. Then, we have $\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}\mathbbm{1}(\tau_{t}\in\mathcal{{H}}_{t}({\bar{\tau}}))\cdot p_{t}(\bar{\tau})=p_{t}(\tau_{t})$, recovering the Exp3 unbiased estimator in (2) from our unbiased estimator (8).
Connection to EW. Our algorithm can be also reduced to EW. Specifically, let $\mathcal{{H}}_{t}({\tau})=\mathcal{{H}}$ for any $\tau$, i.e., full feedback. Then, we have $\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}\mathbbm{1}(\tau_{t}\in\mathcal{{H}}_{t}({\bar{\tau}}))\cdot p_{t}(\bar{\tau})=1$, making our estimator (8) the same as the loss $\ell_{t}$.
@@
Algorithm 1 Exp3 for Online Selective Generation with Partial Feedback Unlocking (ExSUL)
1:procedure ExSUL($T,\mathcal{{H}},\alpha,\lambda,\eta,f,G$)
@@
4:         Observe $\mathbf{x}_{t}$ or equivalently $f_{t}\coloneqq f(\mathbf{x}_{t},G(\mathbf{x}_{t}))$
5:         Predict $\hat{{S}}(\mathbf{x}_{t};\tau_{t})$ where $\tau_{t}\sim p_{t}(\tau)=\sum_{\tau\in\mathcal{{H}}}\delta(\tau)\cdot{w_{t}(\tau)}/{\sum_{\tau\in\mathcal{{H}}}w_{t}(\tau)}$
@@
7:         if  $\hat{{S}}(\mathbf{x}_{t};\tau_{t})\neq\texttt{IDK}$  then
8:              $\mathcal{{H}}_{t}(\tau_{t})\leftarrow\{\tau\in\mathcal{{H}}\mid\tau\leq f_{t}\}$
9:         else
10:              $\mathcal{{H}}_{t}(\tau_{t})\leftarrow\{\tau\in\mathcal{{H}}\mid\tau>f_{t}\}$
@@
[…]
14:procedure ComputeLoss($\hat{\mathbf{s}},b,\alpha,\lambda$)
15:     $a\leftarrow\mathbbm{1}(\hat{\mathbf{s}}=\texttt{IDK})$
16:     $d\leftarrow\mathbbm{1}(\hat{\mathbf{s}}\neq\texttt{IDK})\cdot b-\alpha\,\mathbbm{1}(\hat{\mathbf{s}}\neq\texttt{IDK})+\alpha$
17:     $\ell\leftarrow a+\lambda d$
18:     return $\ell$
4.3 Regret Bound
Our Algorithm 1 has a sublinear regret bound. From this regret bound, we construct the FDR bound by using Lemma 1. See Appendix G.4 for a proof on our novel regret bound.

### v2

@@ 2 Problem: Online Selective Generation with Partial Feedback
We consider online learning of a selective generator under partial feedback for language models. See Appendix B for its preliminary. Let $\mathcal{{W}}$ be a set of tokens and $\mathcal{{X}}=\mathcal{{Y}}\coloneqq\cup_{i=0}^{\infty}\mathcal{{W}}^{i}$ be a set of (token) sequences. Here, given a generator $G:\mathcal{{X}}\to\mathcal{{Y}}$, we consider a time-varying selective generator $\hat{{S}}_{t}:\mathcal{{X}}\to\mathcal{{Y}}\cup\{\texttt{IDK}\}$ that abstains from answering if a generated answer $G(\mathbf{x}_{t})$ at time $t$ is uncertain, i.e., $\hat{{S}}_{t}(\mathbf{x}_{t})\coloneqq\begin{cases}G(\mathbf{x}_{t})&\text{if~}\hat{{s}}(\mathbf{x}_{t},G(\mathbf{x}_{t}))=1\\ \texttt{IDK}&\text{otherwise}\end{cases},$ where $\hat{{s}}:\mathcal{{X}}\times\mathcal{{Y}}\to\{0,1\}$ is a selection function and IDK represents “I don’t know”. Following the conventional selective prediction literature, we consider the scalar parameterization of the selection function $\hat{{s}}$ given a scoring function $f:\mathcal{{X}}\times\mathcal{{Y}}\to{[0,1)}$, i.e., $\hat{{s}}(\mathbf{x}_{t},G(\mathbf{x}_{t}))\coloneqq\mathbbm{1}\left(f(\mathbf{x}_{t},G(\mathbf{x}_{t}))\geq\tau\right),$ thus the selective generator $\hat{{S}}$ is parameterized by $\tau$, denoted by $\hat{{S}}(\cdot;\tau)$. Here, the scoring function $f$ measures the confidence of $G(\mathbf{x}_{t})$ being an answer of $\mathbf{x}_{t}$. Also, we specifically consider that $\tau$ is from the finely-quantized, finite space of $[0,1]$, i.e., $\mathcal{{H}}=\left\{k/(H-1):k=0,1,\dots,H-1\right\}$, where $H$ is the number of hypotheses.
To formalize supervision signals in online learning, we define a feedback function $\mathcal{{E}}_{t}:\mathcal{{H}}\rightarrow\{0,1\}$ that outputs feedback given a selective generator $\hat{{S}}$, i.e., $\mathcal{{E}}_{t}(\hat{{S}})\coloneqq\mathbbm{1}(\hat{S}(\mathbf{x}_{t})\neq_{E}\mathbf{y}_{t})$, and assume that a learner does not observe a ground-truth $\mathbf{y}_{t}$ but instead receives partial feedback $\mathcal{{E}}_{t}(\hat{{S}}_{t})$, which we simply denote as $e_{t}$ in the paper. Here, $A\neq_{E}B$ means that $A$ and $B$ are different in terms of a given correctness relation $E$, e.g., textual-entailment (Bowman et al., 2015); thus, $e_{t}=0$ if the generation is correct and $e_{t}=1$ if it is incorrect or $\hat{S}_{t}(\mathbf{x}_{t})=\texttt{IDK}$ by definition. This setup reflects real-world scenarios, where a user typically provides $e_{t}=0$ or $1$, e.g., thumbs-up or down, instead of $\mathbf{y}_{t}$.
To learn an online selective generator with this partial feedback, we consider online learning in a non-stochastic assumption: each step $t$ until a time horizon $T$, (1) an adversary chooses $\mathbf{x}_{t}\in\mathcal{{X}}$ and $\mathbf{y}_{t}\in\mathcal{{Y}}$, which determine the feedback function $\mathcal{{E}}_{t}$ (2) a learner observes $\mathbf{x}_{t}\in\mathcal{{X}}$ and predicts $\hat{{S}}_{t}(\mathbf{x}_{t})$ where $\hat{{S}}_{t}$ is drawn from a learned distribution $p_{t}$ over selective generators, and (3) the leaner observes partial feedback $e_{t}$ and update $p_{t}$ by using it. Here, we consider an adaptive adversary, i.e., $(\mathbf{x}_{t},\mathbf{y}_{t})$ is drawn from a distribution that depends on the learner’s previous decisions $\hat{{S}}_{1},\dots,\hat{{S}}_{t-1}$ but crucially not on the current decision $\hat{{S}}_{t}$.
Goal. Under this learning setup, our primary goal is to control the false discovery rate (FDR) of a selective generator at a desired level $\alpha\in[0,1]$ up to a time horizon $T$. To this end, we define the FDR risk as follows:
$\mathcal{{R}}_{T}^{\textbf{FDR}}\coloneqq\sum_{t=1}^{T}\Big[\mathbbm{1}(\hat{{S}}_{t}(\mathbf{x}_{t})\neq\texttt{IDK}\wedge e_{t}=1)-\alpha\mathbbm{1}(\hat{{S}}_{t}(\mathbf{x}_{t})\neq\texttt{IDK})\Big].$ (1)
The main objective is to ensure $\mathcal{{R}}_{T}^{\textbf{FDR}}\leq 0$, equivalent to requiring the empirical FDR less than $\alpha$:
$\textbf{FDR}_{T}\coloneqq\frac{\sum_{t=1}^{T}\mathbbm{1}\left(\hat{{S}}_{t}(\mathbf{x}_{t})\neq\texttt{IDK}~\wedge~e_{t}=1\right)}{\sum_{t=1}^{T}\mathbbm{1}\left(\hat{{S}}_{t}(\mathbf{x}_{t})\neq\texttt{IDK}\right)}\leq\alpha,$ (2)
where $\textbf{FDR}_{T}=\alpha$ if $\hat{{S}}_{t}(\mathbf{x}_{t})=\texttt{IDK}$ for all $t$. Then, the goal is to learn a distribution $p_{t}$ over selective generators such that a drawn selective generator $\hat{{S}}_{t}\sim p_{t}$ controls $\mathcal{{R}}_{T}^{\textbf{FDR}}$ close to zero, i.e., $\sum_{t=1}^{T}\big[\mathbbm{1}(\hat{{S}}_{t}(\mathbf{x}_{t})\neq\texttt{IDK}\wedge e_{t}=1)-\alpha\mathbbm{1}(\hat{{S}}_{t}(\mathbf{x}_{t})\neq\texttt{IDK})\big]\leq\varepsilon(T),$ where $\varepsilon(T)$ is some non-increasing function in $T$. Note that the learner controls $\mathcal{{R}}_{T}^{\textbf{FDR}}$ while the adaptive adversary tries to maximize it. Additionally, while FDR controllability is the primary objective, we also wish to minimize selection inefficiency $\textbf{Ineff}_{T}\coloneqq\frac{1}{T}\sum_{t=1}^{T}\mathbbm{1}(\hat{{S}}_{t}(\mathbf{x}_{t})=\texttt{IDK}).$
3 ExSUL: Online Selective Generation with Feedback Unlocking
We leverage a regret perspective from bandit problems to design a learning algorithm for selective generation that controls the FDR under partial feedback. As a preliminary, we first introduce the necessary background on regret minimization in adversarial bandits (Section 3.1). We then reduce online selective generation problems to adversarial bandit problems with partial feedback (Section 3.2). This reduction enables us to leverage any regret minimization algorithms with their regret bounds for selective generation algorithms with FDR bounds. However, the connection between regret and FDR is missing. To fill this gap, we introduce a novel conversion lemma from the regret to the FDR (Section 3.3).
However, under partial feedback, simply leveraging existing bandit algorithms may not lead to sample efficiency without exploiting unique properties of selective generation. To address this, we propose a novel sample-efficient method for online selective generation under partial feedback. In particular, we extend the Exp3-IX algorithm (Neu, 2015) to learn selective generators, exploiting the unique structure of their selection functions by feedback unlocking, theoretically and empirically demonstrating its efficient regret bound (Section 3.4), which eventually leads to an FDR bound by using our conversion lemma. Note that online selective generation under full feedback can be devised in a similar way. See Appendix D.1 for the detail via exponential weighting (EW).
3.1 Regret Minimization and Adversarial Bandits
The adversarial bandit problem (Auer et al., 2002) is formulated as an interactive game between a learner and an adversary. At each round $t\in{1,\dots,T}$, the learner selects an arm $\tau_{t}\in\mathcal{{H}}$ among a set of arms $\mathcal{{H}}$, while the adversary simultaneously chooses a loss function $\ell_{t}:\mathcal{{H}}\to\mathbb{R}_{\geq 0}$. The learner then observes only the incurred loss $\ell_{t}(\tau_{t})$ as partial feedback. The learner’s performance is evaluated by the regret (Bubeck et al., 2012), i.e.,
$\textbf{Reg}_{T}\coloneqq\sum_{t=1}^{T}\ell_{t}(\tau_{t})-\min_{\tau\in\mathcal{{H}}}\sum_{t=1}^{T}\ell_{t}(\tau),$ (3)
and the goal of the learner is to minimize $\textbf{Reg}_{T}$ so that it grows sublinearly in $T$. Here, the adaptive adversary maximizes $\textbf{Reg}_{T}$ by choosing $\ell_{t}$ based on the previously chosen arms by the learner, i.e., $\tau_{1},\dots,\tau_{t-1}$. This makes $\textbf{Reg}_{T}$ a random variable, so we analyze high-probability or expected regret bounds. For further details on regret minimization and Exp3-IX (Neu, 2015), see Appendix B.3 and B.4.
@@ 3.2 Reduction: From Online Selective Generation to Adversarial Bandits
models
selective generators $\mathcal{{H}}$
finite arms $\mathcal{{H}}$
feedback $e_{t}$ $\ell_{t}(\tau_{t},\alpha)$
@@
Here, we map the components of online selective generation with partial feedback to those of adversarial bandits to leverage existing algorithms and regret bounds. In particular, a parameter $\tau\in\mathcal{{H}}$ of our learner corresponds to a bandit arm in adversarial bandits.
For the feedback, we first define a special loss $\ell_{t}(\tau,\alpha)$ and connect this to $e_{t}$. In particular, we introduce the following two key components: $a_{t}(\tau)\coloneqq\mathbbm{1}(\hat{{S}}(\mathbf{x}_{t};\tau)=\texttt{IDK})$, which measures the selection inefficiency at time $t$, called inefficiency loss, and $d_{t}(\tau,\alpha)\coloneqq\mathbbm{1}(\hat{{S}}(\mathbf{x}_{t};\tau)\neq\texttt{IDK}\wedge e_{t}=1)-\alpha\mathbbm{1}(\hat{{S}}(\mathbf{x}_{t};\tau)\neq\texttt{IDK})+\alpha$, which measures the violation of the FDR risk at time $t$ with a margin by $\alpha$ to penalize the IDK response, called FDR loss with a margin. Based on these, we define the loss function for bandits as follows:
$\ell_{t}(\tau,\alpha)\coloneqq\frac{a_{t}(\tau)+\lambda d_{t}(\tau,\alpha)}{1+\lambda}\in\left\{0,\frac{\lambda}{1+\lambda},\frac{1+\lambda\alpha}{1+\lambda}\right\}\in[0,1],$ (4)
where $\lambda$ is a hyperparameter that controls the trade-off between the FDR loss and the efficiency loss. Note that from the bandits’ perspective, the learner receives feedback in two steps: get feedback $e_{t}$ from the adversary and uses this to compute the loss $\ell_{t}(\tau_{t},\alpha)$ for the chosen arm $\tau_{t}$.
Finally, the link between two metrics for online selective generation and adversarial bandits remains to be established. In the following section, we propose a novel conversion lemma to connect $\textbf{FDR}_{T}$ and $\textbf{Ineff}_{T}$ to $\textbf{Reg}_{T}$. See Figure 2 for an illustration of our full learning pipeline from online selective generation to adversarial bandits and back again.
3.3 Regret-to-FDR Conversion
We mainly leverage algorithms that minimize $\textbf{Reg}_{T}$ for learning selective generators. Yet, the key requirement for selective generation is the FDR guarantee of a learner at a desired level. To this end, we introduce a novel perspective on the connection of a regret bound to an FDR bound. In particular, any learner, including learners for bandits, which minimizes the regret (3) with the loss (4), controls the FDR (1). This is achievable as the designed loss (4) penalizes when FDR loss at each $t$ is higher than $\alpha$, weighted by a parameter $\lambda$. See Appendix H.3 for a proof and details on choosing $\lambda$.
Lemma 1.
Let $T\in\mathbb{N}$ and $\alpha\in(0,1)$. For any $(\mathbf{x}_{t},\mathbf{y}_{t})$ sequences, leading to any loss sequences $\ell_{t}$ of (4), we have
$\mathcal{{R}}_{T}^{\textbf{FDR}}\leq\frac{(1-\textbf{Ineff}_{T})+(1+\lambda)\textbf{Reg}_{T}/T}{\lambda}=\frac{(1-\textbf{Ineff}_{T})}{T^{\nicefrac{{1}}{{4}}}}+\frac{(1+T^{\nicefrac{{1}}{{4}}})\textbf{Reg}_{T}}{T^{\nicefrac{{5}}{{4}}}},$ (5)
where the last equality holds if we take $\lambda=1/T^{\nicefrac{{1}}{{4}}}$.
Importantly, this lemma is applicable to learning for both full and partial feedback, as it is agnostic to feedback mechanisms. Also, the lemma implies that if $\textbf{Reg}_{T}$ has a sublinear bound, satisfied by most bandit learners, then $\mathcal{{R}}_{T}^{\textbf{FDR}}$ gets closer to 0 at most the rate of $\mathcal{{O}}(1/T^{1/4})$, i.e., $\mathcal{{R}}_{T}^{\textbf{FDR}}\leq\mathcal{{O}}(1/T^{1/4})$.
The Connection between $\mathcal{{R}}_{T}^{\textbf{FDR}}$ and $\textbf{FDR}_{T}$. We can derive the bound for $\textbf{FDR}_{T}$ by dividing (5) by $1-\textbf{Ineff}_{T}$ if $\textbf{Ineff}_{T}\neq 1$, i.e.,
$\textbf{FDR}_{T}\leq\alpha+\frac{1}{T^{\nicefrac{{1}}{{4}}}}+\frac{(1+T^{\nicefrac{{1}}{{4}}})\textbf{Reg}_{T}}{T^{\nicefrac{{5}}{{4}}}(1-\textbf{Ineff}_{T})},$
which remains sublinear in $T$ unless $\textbf{Ineff}_{T}$ converges to $1$. In other words, $\textbf{FDR}_{T}$ may fail to be bounded if $1-\textbf{Ineff}_{T}\approx 0$ as $\sum_{t=1}^{T}e_{t}\approx T$, due to the nature of the empirical metric with the variance induced by randomized bandit algorithms. For instance, if the generator $G$ consistently produces incorrect outputs, the learner will try to converge to an always-abstaining solution. However, due to randomization, if the learner mistakenly accepts even a single sample in the early stage of learning, then regardless of convergence to an always-abstaining solution, the empirical FDR may still exceed the desired level $\alpha$. We highlight that this reflects an intrinsic limitation of the empirical FDR evaluation in sequential prediction. Thus, we consider the FDR risk, where our adaptive adversary assumes to maximize the $\mathcal{{R}}_{T}^{\textbf{FDR}}$, where it does not consider this pathological case.
Figure 2: An example of our proposed framework for online selective generation. At each step $t$, (1) the user provides an input $\mathbf{x}_{t}$, (2) the learner selects an arm $\tau_{t}\sim p_{t}$, (3-4) selectively generates $\hat{{S}}_{t}(\mathbf{x}_{t};\tau_{t})$ , (5-6) the user provides partial feedback $e_{t}$, (7) the loss $\ell_{t}(\tau_{t})$ is computed from $e_{t}$, and (8) the learner update $p_{t}$ by the loss estimator $\tilde{\ell}_{t}$. Note that the user can be modeled as an adversary.
3.4 ExSUL: Exp3-IX for Online Selective Generation with Feedback Unlocking
@@
As for the adversarial bandit learner, we leverage Exp3-IX (Neu, 2015), which is based on EW (Littlestone & Warmuth, 1994) for partial feedback. The key challenge from full to partial feedback is the slow convergence of learning due to the lack of feedback information. To address this, we exploit the unique loss function structure of selective generators (4) to unlock feedback of other arms by a chosen arm, dubbed by feedback unlocking. In particular, by the monotonicity of a selection function in $\tau$, i.e., $\hat{{s}}(\mathbf{x})\coloneqq\mathbbm{1}(f_{t}\geq\tau)$ where $f_{t}\coloneqq f(\mathbf{x}_{t},G(\mathbf{x}_{t}))$ and the definition of a selective generator $\hat{{S}}$, the following relation among feedback holds, where $e_{t}$ is observed feedback by a chosen arm $\tau_{t}$:
$e_{t}=\mathbbm{1}(\hat{{S}}(\mathbf{x}_{t};\tau_{t})\neq_{E}\mathbf{y}_{t})=\begin{cases}\mathbbm{1}(\hat{{S}}(\mathbf{x}_{t};\tau)\neq_{E}\mathbf{y}_{t})\text{~for~}\tau\leq f_{t}&\text{if~}\hat{{S}}(\mathbf{x}_{t};\tau_{t})\neq\texttt{IDK}\\ \mathbbm{1}(\hat{{S}}(\mathbf{x}_{t};\tau)\neq_{E}\mathbf{y}_{t})\text{~for~}\tau>f_{t}&\text{if~}\hat{{S}}(\mathbf{x}_{t};\tau_{t})=\texttt{IDK}\end{cases}.$ (6)
@@
To exploit this feedback unlocking within Exp3-IX, we propose a novel loss estimator
@@
$\ell_{t}(\tau,\alpha\mid\mathcal{{H}}_{t}({\tau_{t}}))\coloneqq\frac{\ell_{t}(\tau,\alpha)}{\gamma_{t}+\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\bar{\tau}}))\cdot p_{t}(\bar{\tau})}\cdot\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\tau_{t}}))\in[0,\infty),$ (7)
@@
where $\gamma_{t}>0$, $\mathcal{{H}}_{t}({\tau_{t}})\coloneqq\{\tau\in\mathcal{{H}}\mid\tau\leq f_{t}\}$ if $\hat{{S}}(\mathbf{x}_{t};\tau_{t})\neq\texttt{IDK}$ and $\mathcal{{H}}_{t}(\tau_{t})\coloneqq\{\tau\in\mathcal{{H}}\mid\tau>f_{t}\}$ otherwise – see Figure 29 for the visualization of $\mathcal{{H}}_{t}(\tau)$. With learning rate $\eta_{t}>0$, Algorithm 1 presents our fixed-parameter extension of Exp3-IX with partial feedback unlocking. Note that our algorithm can be reduced to Exp3-IX by letting $\mathcal{{H}}_{t}({\tau})=\{\tau_{t}\}$ for any $\tau$, recovering the Exp3-IX estimator in (9) from our estimator (7). Also, our algorithm does not use $f_{t}$ to choose $\tau_{t}$.
@@
Algorithm 1 Exp3-IX for Online Selective Generation with Feedback Unlocking (ExSUL)
1:procedure ExSUL($T,\mathcal{{H}},\alpha,\lambda,\eta,f,G,\gamma$)
@@
4:   Choose $\tau_{t}\sim p_{t}(\tau)=\sum_{\tau\in\mathcal{{H}}}\delta(\tau)\cdot{w_{t}(\tau)}/{\sum_{\tau\in\mathcal{{H}}}w_{t}(\tau)}$
5:   Observe $\mathbf{x}_{t}$
@@
7:   $\tilde{\ell}_{t}(\tau,\alpha)\leftarrow\textsc{Unlocking}(\mathcal{{H}},\alpha,\lambda,\gamma,\tau_{t},\mathbf{x}_{t},e_{t})$
[…]
9:procedure Unlocking($\mathcal{{H}},\alpha,\lambda,\gamma,\tau_{t},\mathbf{x}_{t},e_{t}$)
10:  $f_{t}\leftarrow f(\mathbf{x}_{t},G(\mathbf{x}_{t}))$
11:  $\mathcal{{H}}_{t}(\tau_{t})\leftarrow\begin{cases}\{\tau\in\mathcal{{H}}\mid\tau\leq f_{t}\}&\text{if~}\hat{{S}}(\mathbf{x}_{t};\tau_{t})\neq\texttt{IDK},\\ \{\tau\in\mathcal{{H}}\mid\tau>f_{t}\}&\text{otherwise.}\end{cases}$
@@
14:  return $\ell_{t}\left(\tau,\alpha\mid\mathcal{{H}}_{t}(\tau_{t})\right)$
Regret Bound. Based on the original proof of Exp3-IX but along with our novel loss estimator (7), our Algorithm 1 achieves a sublinear regret bound, which leads to faster convergence of the FDR guarantee by Lemma 1. See Appendix H.4 for a general proof on our regret bound.